what is the melting point of a solution in which 3 grams of sodium chloride is added to 170 ml of water? The cryoscopic constant of water is Kf=1.86Kkgmol-1.
Answer #1

we have to use the formula

delta T = m x kf x i

m = molality we need to calculate

moles of NaCl = 3 / 58.45 = 0.0513

molality = 0.0513 / 0.17 Kg water

= 0.3019 molal

NaCl can dissociate in to Na+ + Cl- so want haf facter i = 2

put all the values in above equation

delta T = 0.3019 x 1.86 x 2

delta T = 1.1231

1.1231 lesser than zero that means

T = 0-1.1231

= -1.1231 ºC